### v1

# ContPhy: Continuum Physical Concept Learning and Reasoning from Videos

Abstract
We introduce the Continuum Physical Dataset (ContPhy), a novel benchmark for assessing machine physical commonsense. ContPhy complements existing physical reasoning benchmarks by encompassing the inference of diverse physical properties, such as mass and density, across various scenarios and predicting corresponding dynamics. We evaluated a range of AI models and found that they still struggle to achieve satisfactory performance on ContPhy, which shows that current AI models still lack physical commonsense for the continuum, especially soft-bodies, and illustrates the value of the proposed dataset. We also introduce an oracle model (ContPRO) that marries the particle-based physical dynamic models with the recent large language models, which enjoy the advantages of both models, precise dynamic predictions, and interpretable reasoning. ContPhy aims to spur progress in perception and reasoning within diverse physical settings, narrowing the divide between human and machine intelligence in understanding the physical world. Project page: https://physical-reasoning-project.github.io.
1 Introduction
Humans are capable of comprehending the physical properties of various substances, including rigid objects and soft objects, understanding their dynamic interactions in complex environments, and predicting their corresponding dynamic changes. In fact, this innate ability to understand and reason about the physical world plays a crucial role in shaping our understanding of nature and the development of scientific knowledge (Kill & Kim, 2020).
As depicted in Fig. 1, objects like solids and liquids in nature often exhibit different properties, and these objects of different properties couple together to build our complex physical world. As humans, we are able to distinguish objects’ physical properties by observing their interactions. We know that the clear liquid in Fig. 1 (a) at the bottom has a higher density than the yellow liquid on the top; we know that the dynamic pulley in Fig. 1 (c) could help us to pull the cargo up more easily. These innate human skills raise an intriguing question: can current AI models have the physical common sense to infer physical properties of the continuum111Continuum encompasses various bodies like liquids, soft materials (e.g., ropes), rigid bodies and articulated bodies (e.g., pulleys). and predict corresponding dynamics?
Recently, a series of benchmarks (Riochet et al., 2018; Rajani et al., 2020; Bear et al., ), have been developed to study machine models’ effectiveness for physical reasoning. However, there have been limitations that make them non-ideal for the assessment of whether machine models have human-like physical reasoning abilities. Firstly, most benchmarks mainly deal with simple visual primitives like spheres, cubes, and collision events of rigid objects only. It remains doubtful whether the conclusions based on these simple scenes will still hold in more comprehensive visual scenarios with the coupling of soft objects and their interaction with rigid objects. There have also been benchmarks like Physion (Bear et al., ) that were developed to evaluate machine models’ physical reasoning abilities in different scenarios. However, objects in Physion are of the same physical parameters without any variance (e.g. solids with the same mass and water with the same density). Moreover, Physion only requires models to predict whether two objects will come into contact after the observed video ends. It has not incorporated natural language to answer other challenging questions like predicting dynamics in counterfactual scenes and selecting actions to achieve a goal.
To this end, we aim to build a Continuum Physical Dataset (ContPhy) to thoroughly evaluate and diagnose machine models’ physical reasoning performance in comprehensive physical environments. The design of ContPhy aims to achieve two goals: 1) covering diverse physical scenarios and 2) supporting comprehensive natural language tasks. To achieve the first goal, we adopt the physical engine (Haas, 2014) to simulate diverse videos with dense supervision signals. As shown in Fig. 2, the simulated physical scenes include scenes with the coupling of different liquids, deformable cloths, pulley systems, and elastoplastic balls. Another goal of the built dataset is to propose diverse physical reasoning tasks in the form of video question answering. We achieve this goal with a carefully designed question engine. The question engine takes the dense simulated video annotation as input and generates different questions based on pre-defined textual templates. Sample questions can be found in Fig. 2. It asks challenging questions such as “If the red stick were removed, would most orange fluid flow into the cyan container?” and “Is the mass of the sphere greater than half that of the red cube?”, which requires the model to have a deep understanding of physical scenes and reason about their dynamics.
We also evaluate a series of traditional AI models (Hudson & Manning, 2018; Li et al., 2022a; Le et al., 2020) and recent multimodal large language models (Team et al., 2023; Achiam et al., 2023) on ContPhy. We found that the performance of these models is far from satisfactory, demonstrating the proposed ContPhy benchmark’s value and indicating the necessity of more advanced models with better physical common sense.
To better investigate the characteristics of ContPhy and show insights to build stronger physical reasoning models, we introduce an oracle model, ContPRO that marries two powerful research ideas, particle-based models (Li et al., 2019; Sulsky et al., 1995) for dynamic predictions and the recent large language models (Ouyang et al., 2022) for complex language reasoning. While it requires more supervision signals (i.e. particle-based representation for the scenes), ContPRO achieves the best overall performance.
To summarize, the contribution of the paper lies in three aspects. First, we introduce a pioneering benchmark for physical reasoning that encapsulates a wide spectrum of physical properties such as mass, density, elasticity, and deformability. Complementing this, we have developed a meticulously crafted question engine capable of synthesizing a variety of complex physical reasoning queries. Second, we extensively evaluate the proposed benchmark with multiple machine models to study the characteristics and show insights into physical reasoning model development. Finally, we develop an oracle model for the benchmark, which combines symbolic representation with particle-based dynamic models for physical understanding and reasoning.
Dataset | Question | Rationales | Diverse | Goal-driven | Interaction | Counterfactual |
Answering | Scenarios | Questions | of soft objects | Property Dynamics | ||
IntPhys (Riochet et al., 2018) | ||||||
ESPRIT (Rajani et al., 2020) | ||||||
Cater (Girdhar & Ramanan, 2020) | ||||||
CoPhy(Baradel et al., 2020) | ||||||
CRAFT (Ates et al., 2020) | ||||||
CLEVRER (Yi et al., 2020) | ||||||
Physion (Bear et al., ) | ||||||
ComPhy (Chen et al., 2022) | ||||||
CRIPP-VQA (Patel et al., 2022) | ||||||
ContPhy (Ours) |
2 Related Work
Physical Reasoning. Our work is closely related to Physical Reasoning benchmarks (Rajani et al., 2020; Girdhar & Ramanan, 2020; Baradel et al., 2020; Bear et al., ; Li et al., 2022b, c). We summarize the key features of these various benchmarks and compare them against our benchmark in table 1. Early benchmarks (Riochet et al., 2018; Rajani et al., 2020) simulate physical scenes with visual primitives and test models’ physical intuition. Later, CLEVER (Yi et al., 2020), ComPhy (Chen et al., 2022), and CRIPP-VQA (Patel et al., 2022) extend the simple visual primitives with natural language and asked questions about rigid bodies’ collisions. Recently, Physion (Bear et al., ; Tung et al., 2023) provides more complex visual scenes and requires models to predict whether two objects will come into contact in future frames. As summarized in table 1, the proposed ContPhy is the only benchmark that contains soft objects with different physical parameters and asks diverse language-based questions about dynamics in counterfactual and goal-planning scenarios.
Video Question Answering. Our paper is also related to Video Question Answering (VQA) (Lei et al., 2018; Zadeh et al., 2019; Jang et al., 2017; Chen et al., 2021; Wu et al., 2021; Ding et al., 2021), which mainly requires machine models to answer questions about a given image or video’s content like visual attributes, actions, activity, and social events. However, existing VQA datasets (Zadeh et al., 2019; Xu et al., 2016; Wang et al., 2019) still typically assess abilities in visual perception, recognizing objects, shapes, and colors, and understanding human-centric actions. In this paper, we aim to build a benchmark that evaluates AI models’ comprehensive physical reasoning abilities.
Physical Benchmarks for Soft Bodies. Recently, there has been growing interest in the properties and dynamics of soft-bodied objects (Xiang et al., 2020; Gan et al., 2020; Macklin et al., 2014; Xian et al., 2023; Haas, 2014). Much of the research has concentrated on creating simulations of deformable objects and fluids to advance robotic manipulation and cognitive experimentation. Leveraging simulation tools, we can simulate deformable objects and fluids with varying physical parameters, enabling collaboration with natural language for physical commonsense reasoning.
3 Dataset
The proposed ContPhy dataset aims to assess the reasoning abilities of AI models across a wide spectrum of physical scenes of the continuum encompassing rigid bodies, soft bodies, and fluids, with massive physical properties treated as variables. In this section, we outline the dataset construction process. In Section 3.1, we describe the dataset’s diversity in physical scenes and provide an introduction to each scenario. Section 3.2 explains the well-structured nature of our question dataset, encompassing properties and dynamics. In Section 3.3, we elucidate how a physical engine is employed to simulate diverse scenes with varying properties, the development of a question engine for question generation, and steps taken to mitigate dataset bias through statistical analysis.
3.1 Diverse Physical Scenarios
We design four physical scenarios to visually illustrate various continuum physical dynamics and to study different physical behaviors across different object materials with varying physical properties.
Diversity. The diversity of our video data arises from a wide range of materials with varying properties and changing dynamic phenomena. We include these materials and their physical properties in our scenarios compositionally, requiring models to obtain a deeper understanding of these scenes. First, this encompasses rigid bodies, ropes, cloths, plastoelastic balls, fluids, and articulated bodies, as well as their couplings, expanding our dataset diversity. For instance, Fig 1 (c) portrays the coupling of deformable ropes and articulated pulleys. Second, a key feature that distinguishes the proposed ContPhy dataset from existing benchmarks like CLEVRER (Yi et al., 2020) and Physion (Bear et al., ), is the inclusion of varying physical properties and commonsense concepts that are generally acknowledged by common people. These properties could scarcely be inferred through static images, involving mass, density, tension, friction, stretchiness, bending stiffness, elasticity, and plasticity. Such variation can alter the dynamics and generate differing future states, enriching the dataset with a wider range of observable physical behaviors. As an example, the liquid’s position in Fig. 1 (a) is jointly determined by its density and its interaction with the nearby liquids.
Scenarios. We design four physical dynamic scenarios to comprehensively benchmark models’ cognitive ability on the continuum. a) Liquid Dynamics: In Fig. 2 A, we present a liquid hourglass-like device. Different liquids, each with unique colors and densities, are released from upper emitters, flowing through fixed sticks, altering directions, and finally reaching containers at the bottom. This setup reveals distinct behaviors arising from interactions between fluids with varied densities and dynamic trajectories. Our research focuses on investigating the physical properties and trajectories of these liquids. b) Cloths Manipulation. As shown in Fig. 2 B, two cloth pieces with distinct stretching, bending, and frictional characteristics are pulled over objects, inducing potential collision events. The released fabric obstructs object views but outlines their shapes through deformations. Objects may topple if they surpass a height threshold or have low mass. This test assesses models’ ability to discern fabric properties and predict spatial behaviors of concealed objects based on the dynamic 3D surface geometry of the fabric. c) Rope Pulley System: In Fig. 2 C, a wall-mounted arrangement of pulleys, both movable and fixed and anchor points are depicted. Objects with varying masses interact, resulting in diverse motion patterns. The model’s main goal is to identify tension distributions within this basic rope system. It must also recognize correlations or constraints among moving objects, such as coordinated loads and pulley rotations on a single rope. Additionally, the model is expected to infer numerical relationships between loads’ masses and rope segment tensions. d) Soft Ball Dynamics: Fig. 2 D illustrates a playground with colored obstacles and randomly placed pits. Plastoelastic balls with different deformation resistance and yield stress are launched from varying positions, undergoing dynamic movements, including bouncing and permanent deformation. Some balls may collide with obstacles and fall into pits. This experiment assesses the model’s ability to accurately discern the elasticity and plasticity properties of soft bodies and predict their dynamic behavior.
3.2 Diverse Structured Questions
We categorize questions into two major groups: Physical Property Questions and Dynamics Questions. Fig. 2 shows the question types of the four scenarios. Sample templates are provided in the supplementary material.
Physical Property Questions. We formulated a set of physical property questions across four distinct scenarios. We pose questions about our chosen physical properties which can only be answered by observing object dynamics and interactions, neither static images nor single object behaviors. These questions can be answered with a brief phrase. Models are expected to deduce physical properties based on input video data, which requires physical common sense. Besides, we also inquire about the visible physical properties of objects, such as colors, shapes, and existences, which are shown in static frames.
Dynamics Questions. Dynamic questions can be further categorized into three types: counterfactual, goal-driven, and predictive, respectively concerning potential outcomes of changed conditions, strategies for specific objectives, and predictions about the future. In the fluid and ball scenarios, we crafted questions covering all three types, anticipating models to develop a comprehensive understanding of these scenarios through diverse question templates. For rope and cloth scenarios, we selectively assess a subset of dynamic question types due to scenario complexity. In the rope scenario, only counterfactual and goal-driven questions are included. In the cloth scenario, exclusively predictive questions prompt the model to anticipate outcomes not directly visible under the cloth cover. To increase cognitive challenge, we’ve designed multiple-choice questions with more than two but fewer than five answer choices, requiring models to provide a binary prediction for each option.
3.3 Generation Setup and Statistics
Video Generation. We used the Unity engine (Haas, 2014), an efficient platform, to simulate and render videos. We follow a bottom-up approach to generate videos and their annotation, involving the following sequential steps: a) Sampling: randomly select scene layouts, camera parameters, and initial conditions to create a diverse set of scenarios. b) Initialization: place and configure objects within the scene. c) Pre-simulation: conduct a preliminary simulation to evaluate whether the obtained simulation results align with the expected data distribution. d) Rendering: generate high-quality videos by configured cameras. e) Post-simulation: carry out multiple simulations under varying conditions and record the simulation outputs. f) Output: produce rich sensor data and annotation information, encompassing original video, segmentation, bounding boxes, particles, meshes, collision events, configurations, and other simulation raw data required for question generation. Supplementary material provides more details.
Question Generation. We develop a question engine to generate question-answering pairs step by step. a) Template Design: create massive question and option templates. b) Sampling: retrieve the simulation results, combine the properties of possible objects with predefined templates, sample questions, and options accordingly, and determine correct answers. Target objects possess unique names described by visual attributes like color, shape, orientation, and mobility. c) Re-Sampling: ensure a balanced distribution of answers among the options to prevent answer bias.
Question Statistics. The video content prompted the formulation of numerous questions, with each video featuring one property question and two dynamics questions, except for the rope scenario, which included two property-related questions and two dynamics questions. A total of 2,000 questions were generated for the rope scenario, and 1,500 questions were created for other scenarios. The dataset encompasses 6,500 questions derived from 2,000 videos. We divided the dataset into three subsets: 50% for training, 20% for validation, and 30% for testing. Across the entire dataset, 20% of questions are counterfactual, 11% are goal-driven, 22% are predictive, and the remaining 46% are related to various physical property questions. Further details on the distribution of each question type and templates for each scenario can be found in the supplementary material.
4 ContPRO
Inspired by prior work (Yi et al., 2018, 2020), we propose an oracle neural-symbolic framework named ContPRO for ContPhy. As shown in Fig. 3, we decompose the question-answering task into four main modules, video perception, physical simulation, program parser, and symbolic execution. Given a raw video, the video perception module detects the objects and their associated static attributes with MASK-RCNN detector (He et al., 2017). The physical simulator takes point clouds as input and predicts objects’ dynamics in different scenarios with dynamic prediction models (Jiang et al., 2015; Li et al., 2018). The program parser translates the question query into executable programs with a large language model (LLM). Based on the object attributes and dynamics, the symbolic executor executes the programs to get the answer to the question.
Compared with previous models, ContPRO is the first model that marries LLMs with particle-based physical dynamic models, which eliminates the need for handcraft design and in-domain training of the program parser and enjoys precise dynamic prediction. As it requires point clouds as input, we call it an oracle model.
Video Perception. The video perception module is supported by a MASK R-CNN (He et al., 2017) to densely detect objects’ location in each frame and associated static attributes like color and material. We take the ResNet-50 (He et al., 2016) as the backbone and fine-tune the network with data from the training set of all four scenarios until converges.
Physical Simulation. We choose DPI-Net (Li et al., 2018) for dynamic prediction for rope and cloth as it has shown reasonable dynamic prediction abilities in different materials (Chen et al., 2022; Bear et al., ). We also observe that the Material Point Method (MPM) (Sulsky et al., 1995) can estimate physical properties and dynamics better for fluids and objects of varying plasticity with its differentiable mechanism during inference. Thus, we adopt MPM to predict dynamics for scenes of fluid and soft balls.
For DPI-Net, we train it with the cloth and rope data from the training set. For the MPM model, we first initialize the physical properties of the object with a fixed value and gradually optimize its value with gradient descent. We set the optimized loss target to be the chamfer loss between the groups of 3D points and the predicted points.
Large Language Model as Program Parser. Traditional neuro-symbolic models (Andreas et al., 2016; Yi et al., 2018) usually train a domain-specific sequence-to-sequence model to translate the natural language query into executable programs, which requires manual implementation of each symbolic operator and show problems in generalizing to questions out of the training set distribution. Motivated by the recent ViperGPT (Surís et al., 2023), we utilize the large language model, ChatGPT (Ouyang et al., 2022) as the language parser. We further develop a set of dynamics modules and visual perception modules serving as APIs for solution generation. With the provided API access and a pre-defined physical reasoning prompt, we leverage ChatGPT to generate Python code that can be directly executed and interpreted. This module bridges the gap between language comprehension and physical concept understanding.
Program Execution. After extracting objects’ static attributes, physical properties, and dynamic trajectories and parsing the natural language query into an executable program, we execute the program with object states as input and output the predicted answer. We provide more model details in the supplementary material.
5 Experiments
Subset | Settings | Blind Models | Visual Models | MLLMs | ContPRO | Human | ||||||
RND | FRQ | B-LSTM | C-LSTM | MAC | HCRN | ALPRO | Gemini | GPT-4V | ||||
Rope | Prop. | 30.0 | 53.3 | 54.7 | 52.7 | 53.3 | 51.7 | 60.7 | 35.5 | 48.0 | 71.4 | 84.7 |
C-Opt. | 51.3 | 51.6 | 74.0 | 74.0 | 74.2 | 74.3 | 76.2 | 48.2 | 42.0 | 75.6 | 90.2 | |
C-Ques. | 14.7 | 19.0 | 46.0 | 45.0 | 39.8 | 48.1 | 50.7 | 12.0 | 11.3 | 48.8 | 75.0 | |
G-Opt. | 55.2 | 49.7 | 47.4 | 51.2 | 50.3 | 56.0 | 46.2 | 51.6 | 57.0 | 55.8 | 91.9 | |
G-Ques. | 4.5 | 11.2 | 7.9 | 6.7 | 6.7 | 2.3 | 1.1 | 10.3 | 12.1 | 2.3 | 84.0 | |
Fluid | Prop. | 33.3 | 52.7 | 49.3 | 54.0 | 30.0 | 52.7 | 48.0 | 10.0 | 25.0 | 78.0 | 75.8 |
C-Opt. | 52.9 | 57.9 | 56.1 | 55.0 | 56.5 | 52.6 | 56.8 | 47.3 | 53.3 | 75.7 | 82.5 | |
C-Ques. | 6.0 | 17.2 | 7.8 | 8.6 | 6.9 | 4.3 | 6.0 | 5.1 | 5.1 | 36.2 | 60.6 | |
G-Opt. | 59.9 | 63.1 | 57.3 | 57.3 | 51.2 | 67.7 | 62.7 | 44.4 | 53.8 | 77.3 | 75.0 | |
G-Ques. | 7.5 | 36.3 | 22.5 | 22.5 | 17.5 | 41.3 | 32.5 | 11.3 | 7.5 | 60.0 | 64.3 | |
P-Opt. | 53.8 | 50.1 | 51.4 | 51.4 | 53.5 | 50.6 | 53.8 | 52.4 | 50.0 | 90.1 | 73.9 | |
P-Ques. | 4.8 | 12.5 | 12.5 | 12.5 | 12.5 | 1.9 | 12.7 | 5.8 | 13.0 | 68.3 | 42.9 | |
Cloth | Prop. | 46.7 | 41.3 | 56.7 | 46.7 | 59.3 | 52.0 | 48.0 | 42.0 | 49.0 | 60.0 | 81.4 |
P-Opt. | 52.2 | 61.7 | 55.2 | 67.5 | 57.9 | 62.0 | 68.8 | 50.1 | 53.0 | 64.7 | 79.6 | |
P-Ques. | 46.0 | 56.7 | 42.3 | 57.3 | 50.7 | 56.3 | 57.3 | 43.0 | 47.5 | 60.0 | 77.3 | |
Ball | Prop. | 53.5 | 52.0 | 45.3 | 54.7 | 48.0 | 43.3 | 48.0 | 54.0 | 45.0 | 54.0 | 76.9 |
C-Opt. | 53.6 | 65.8 | 66.7 | 64.2 | 66.1 | 65.3 | 63.9 | 60.9 | 66.7 | 71.6 | 93.9 | |
C-Ques. | 30.4 | 48.7 | 43.4 | 41.8 | 3.3 | 28.7 | 40.2 | 29.6 | 46.9 | 57.4 | 90.9 | |
G-Opt. | 55.9 | 52.1 | 53.3 | 54.1 | 58.1 | 57.0 | 56.3 | 54.1 | 51.4 | 68.1 | 89.7 | |
G-Ques. | 30.2 | 38.5 | 16.7 | 20.0 | 18.9 | 38.9 | 4.4 | 24.6 | 18.0 | 52.2 | 84.6 | |
P-Opt. | 50.6 | 67.8 | 68.9 | 67.4 | 64.4 | 61.7 | 65.2 | 51.7 | 45.4 | 92.4 | 72.5 | |
P-Ques. | 25.9 | 51.7 | 45.5 | 45.5 | 46.6 | 1.1 | 3.4 | 25.9 | 17.2 | 88.6 | 58.8 |
In this section, we evaluate the ContPhy dataset. We first introduce the experimental setup and then analyze different models’ performance.
5.1 Experimental Setup
For simplicity, each physical property question is regarded as a classification task among all possible answers. Each dynamic question is treated as a binary classification task for each question-choice pair. For dynamic questions, we report the accuracy for each option and per question. A question is correct only if all choices in this multiple-choice question are correctly answered.
Blind Models. This family of models includes baselines that only rely on question input, to analyze language biases in ContPhy. RND chooses at random a possible answer, or randomly selects between true-false binary answer pairs for every multiple-choice question. FRQ selects the most frequent answer based on the question type. B-LSTM utilizes an LSTM (Hochreiter & Schmidhuber, 1997) to encode the questions only and predict answers.
Visual Models. These models incorporate both visual and language representations for answering questions. C-LSTM extracts video features via ResNet-50 convolutional neural network (CNN) (He et al., 2016) on 25 sampled frames of videos and averages them over time as the visual input. We concatenate this visual input with the question embedding from the last hidden state of LSTM to predict answers. HCRN (Le et al., 2020) uses conditional relational networks to learn relations hierarchically in the video, as well as the questions. MAC (Hudson & Manning, 2018) has competitive results on previous datasets, which uses co-attention mechanism to model both textual and visual information. ALPRO (Li et al., 2022a) is a popular model pre-trained on video-text corpus and achieved state-of-the-art results on several video-language datasets. We fine-tune ALPRO on our dataset based on the official pre-trained checkpoint.
5.2 Evaluation of Physical Reasoning
We summarize the performance of all baselines in Table 2. The results show that different models exhibit distinct performance variances across different question types and scenarios. This indicates that ContPhy can evaluate models’ physical reasoning capabilities in different dimensions.
Performance of Blind Models. Blind models operate and respond solely to textual data, reflecting the quality and structure of question design. Generally, these models have weaker performance than other families of models, showing the importance of cooperating visual information to handle the questions in ContPhy. We also observe that blind models perform similarly to other model families in some scenarios like the goal-driven questions of the rope scenario. We think the reasons are that these dynamic questions are too challenging for current machine models, which require the understanding of physical commonsense and predict information that is not directly observable. Note that human beings can still easily achieve high performance.
Performance of Visual Models. Visual models, which integrate both visual and language representations, exhibit relatively consistent performance across various questions. Among all dynamic questions, they excel on the rope’s counterfactual and cloth’s predictive, but fall short on goal-driven questions. The inherent complexity of goal-driven questions, which require reverse reasoning based on the goal, may account for its bad performance. Among different scenarios and visual models, ALPRO distinguishes itself by its robust overall performance, notably in cloth and rope, which shows the advantages of large-scale video-text pre-training and alignment, emphasizing its effectiveness in complex visual reasoning. Despite these advancements, no visual model has yet achieved top accuracy in all scenarios, underscoring the challenge and significance of our ContPhy.
Performance of Multimodal Large Language Models. Compared with other baselines, both foundation models, Gemini and GPT4-V, fall short in cloth and fluid questions in both property and dynamics levels, showing that they fail to perceive highly deformable objects. However, they perform the best in certain question settings of rope and ball scenarios. Note that these models have never been trained on ContPhy and their visibility is limited to discrete frames in our setup. We think the reason is that objects in rope and ball, e.g. cubes and spheres, are more common to foundation models than those in fluid and cloth, e.g. different colored liquids and clothes. For example, GPT-4V shows its capabilities in counting objects and perceiving object colors, which probably raises the rope property evaluation score. MLLMs also distinguish themselves by their complex reasoning capabilities, accounting for high accuracy in rope’s goal-driven questions. We argue that current foundation models lack the necessary capability to infer physical properties and predict the dynamics of complex scenes with soft objects and fluid.
Performance of ContPRO. ContPRO excels significantly over other machine models on most questions, particularly on property inference and predictive questions. Notably, its accuracy on predictive questions even surpasses human performance. A possible explanation is that we adopt MPM for fluid and ball scenarios, which can precisely predict objects’ positions in the short term, while humans tend to give vague estimations. Such an approach also accounts for its superior performance on the fluid’s goal-driven questions. In addition, ContPRO demonstrates exceptional performance on fluid property inference, achieving an accuracy of 96.9% on “stick number” and 63.5% on “density”. We believe the reason is that we have utilized a fine-tuned Mask R-CNN predictor to identify sticks in videos. For rope and ball scenarios, we employ DPI-Net, a GNN-based simulator, which cannot exhibit absolute advantages over visual models.
Human Performance. We randomly sampled some video-question pairs from the test set to assess the human ability to comprehend the physical properties and dynamic events presented in both video and textual descriptions. To evaluate human performance on ContPhy, 16 people participated in the study. Participants were required to have fundamental English reading skills and a basic physical knowledge background. First, each participant was asked to select a scenario randomly, after which they were presented with distinct video-question pairs. Participants were instructed to answer with a phrase when presented with physical property questions, while for dynamics questions they were required to provide a binary true-false response from available choices. We obtained 460 valid human answers encompassing all scenarios and question types within ContPhy. We can observe from Table 2 that it beats visual models and foundation models in all scenarios. This shows the fundamental ability and strength of humans to perform visual reasoning and inference from videos.
Evaluation Conclusion. The strong human results demonstrate that humans maintain a strong capacity to comprehend both videos and questions, make physical property inferences from given videos, and predict and reason counterfactual hypotheses concerning unseen information. Machine model results show that even state-of-the-art models struggle with answering these physical questions. This indicates that our dataset poses a significant challenge for vision-language models to achieve similar basic physical video understanding ability with human beings. We also propose an oracle model to demonstrate the potential to combine recent large language models with traditional particle-based dynamic simulation for effective physical reasoning.
6 Conclusion
We introduced the Continuum Physical Dataset (ContPhy), a pioneering benchmark for assessing machine models in physical reasoning of the continuum, especially for soft bodies and fluids. This benchmark broadens the scope by covering various physical property inferences for soft bodies across dynamic contexts and predicting their dynamics. Our dataset has enabled the development of AI models with human-like reasoning abilities, comprehending both visual attributes and complex physical properties of objects while solving problems. Despite progress, our evaluation of AI models revealed an ongoing challenge: they struggle to perform well on our benchmark, highlighting their limited physical commonsense for the continuum, especially soft bodies, and fluids. We foresee the ContPhy driving progress in AI perception and reasoning, bridging the gap between human and machine intelligence in the physical world.

### v2

@@
 In fact, this innate ability to understand and reason about the physical world plays a crucial role in shaping our understanding of nature and the development of scientific knowledge (Kill & Kim, 2020).
-As depicted in Fig.
-1, objects like solids and liquids in nature often exhibit different properties, and these objects of different properties couple together to build our complex physical world.
-As humans, we are able to distinguish objects’ physical properties by observing their interactions.
-We know that the clear liquid in Fig.
-1 (a) at the bottom has a higher density than the yellow liquid on the top; we know that the dynamic pulley in Fig.
-1 (c) could help us to pull the cargo up more easily.
+As depicted in Figure 1, objects like solids and liquids in nature often exhibit different properties, and these objects of different properties couple together to build our complex physical world.
+As humans, we are able to distinguish objects’ physical properties by observing their interactions.We know that the clear liquid in Figure 1 (a) at the bottom has a higher density than the yellow liquid on the top; we know that the dynamic pulley in Figure 1 (c) could help us to pull the cargo up more easily.
 These innate human skills raise an intriguing question: can current AI models have the physical common sense to infer physical properties of the continuum111Continuum encompasses various bodies like liquids, soft materials (e.g., ropes), rigid bodies and articulated bodies (e.g., pulleys).
+More details about the physical concept of the continuum can be found in Section 9 in the Appendix.
 and predict corresponding dynamics?
-Recently, a series of benchmarks (Riochet et al., 2018; Rajani et al., 2020; Bear et al., ), have been developed to study machine models’ effectiveness for physical reasoning.
+Recently, a series of benchmarks (Riochet et al., 2018; Rajani et al., 2020; Bear et al., 2021), have been developed to study machine models’ effectiveness for physical reasoning.
 However, there have been limitations that make them non-ideal for the assessment of whether machine models have human-like physical reasoning abilities.
@@
 It remains doubtful whether the conclusions based on these simple scenes will still hold in more comprehensive visual scenarios with the coupling of soft objects and their interaction with rigid objects.
-There have also been benchmarks like Physion (Bear et al., ) that were developed to evaluate machine models’ physical reasoning abilities in different scenarios.
+There have also been benchmarks like Physion (Bear et al., 2021) that were developed to evaluate machine models’ physical reasoning abilities in different scenarios.
 However, objects in Physion are of the same physical parameters without any variance (e.g.
@@
 To achieve the first goal, we adopt the physical engine (Haas, 2014) to simulate diverse videos with dense supervision signals.
-As shown in Fig.
-2, the simulated physical scenes include scenes with the coupling of different liquids, deformable cloths, pulley systems, and elastoplastic balls.
+As shown in Figure 2, the simulated physical scenes include scenes with the coupling of different liquids, deformable cloths, pulley systems, and elastoplastic balls.
 Another goal of the built dataset is to propose diverse physical reasoning tasks in the form of video question answering.
@@
 The question engine takes the dense simulated video annotation as input and generates different questions based on pre-defined textual templates.
-Sample questions can be found in Fig.
-2.
+Sample questions can be found in Figure 2.
 It asks challenging questions such as “If the red stick were removed, would most orange fluid flow into the cyan container?” and “Is the mass of the sphere greater than half that of the red cube?”, which requires the model to have a deep understanding of physical scenes and reason about their dynamics.
@@
 Finally, we develop an oracle model for the benchmark, which combines symbolic representation with particle-based dynamic models for physical understanding and reasoning.
-Dataset | Question | Rationales | Diverse | Goal-driven | Interaction | Counterfactual |
-Answering | Scenarios | Questions | of soft objects | Property Dynamics | ||
-IntPhys (Riochet et al., 2018) | ||||||
-ESPRIT (Rajani et al., 2020) | ||||||
-Cater (Girdhar & Ramanan, 2020) | ||||||
-CoPhy(Baradel et al., 2020) | ||||||
-CRAFT (Ates et al., 2020) | ||||||
-CLEVRER (Yi et al., 2020) | ||||||
-Physion (Bear et al., ) | ||||||
-ComPhy (Chen et al., 2022) | ||||||
-CRIPP-VQA (Patel et al., 2022) | ||||||
+| Dataset | Question | Rationales | Diverse | Goal-driven | Interaction | Counterfactual | |
+| Answering | Scenarios | Questions | of soft objects | Property Dynamics | |||
+| IntPhys (Riochet et al., 2018) | |||||||
+| ESPRIT (Rajani et al., 2020) | |||||||
+| Cater (Girdhar & Ramanan, 2020) | |||||||
+| CoPhy(Baradel et al., 2020) | |||||||
+| CRAFT (Ates et al., 2020) | |||||||
+| CLEVRER (Yi et al., 2020) | |||||||
+| Physion (Bear et al., 2021) | |||||||
+| ComPhy (Chen et al., 2022) | |||||||
+| ACQUIRED (Wu et al., 2023) | |||||||
+| CRIPP-VQA (Patel et al., 2022) | |||||||
 | ContPhy (Ours) |
@@
 Physical Reasoning.
-Our work is closely related to Physical Reasoning benchmarks (Rajani et al., 2020; Girdhar & Ramanan, 2020; Baradel et al., 2020; Bear et al., ; Li et al., 2022b, c).
+Our work is closely related to Physical Reasoning benchmarks (Rajani et al., 2020; Girdhar & Ramanan, 2020; Baradel et al., 2020; Bear et al., 2021; Li et al., 2022b, c).
 We summarize the key features of these various benchmarks and compare them against our benchmark in table 1.
@@
 Later, CLEVER (Yi et al., 2020), ComPhy (Chen et al., 2022), and CRIPP-VQA (Patel et al., 2022) extend the simple visual primitives with natural language and asked questions about rigid bodies’ collisions.
-Recently, Physion (Bear et al., ; Tung et al., 2023) provides more complex visual scenes and requires models to predict whether two objects will come into contact in future frames.
+Recently, Physion (Bear et al., 2021; Tung et al., 2023) provides more complex visual scenes and requires models to predict whether two objects will come into contact in future frames.
 As summarized in table 1, the proposed ContPhy is the only benchmark that contains soft objects with different physical parameters and asks diverse language-based questions about dynamics in counterfactual and goal-planning scenarios.
 Video Question Answering.
-Our paper is also related to Video Question Answering (VQA) (Lei et al., 2018; Zadeh et al., 2019; Jang et al., 2017; Chen et al., 2021; Wu et al., 2021; Ding et al., 2021), which mainly requires machine models to answer questions about a given image or video’s content like visual attributes, actions, activity, and social events.
-However, existing VQA datasets (Zadeh et al., 2019; Xu et al., 2016; Wang et al., 2019) still typically assess abilities in visual perception, recognizing objects, shapes, and colors, and understanding human-centric actions.
+Our paper is also related to Visual Question Answering (VQA) (Lei et al., 2018; Zadeh et al., 2019; Jang et al., 2017; Chen et al., 2021; Wu et al., 2021; Ding et al., 2021; Hong et al., 2023; Chen et al., 2024b, a), which mainly requires machine models to answer questions about a given image or video’s content like visual attributes, actions, activity, and social events.
+However, existing VQA datasets (Zadeh et al., 2019; Xu et al., 2016; Wang et al., 2019; Wu et al., 2023; Wang et al., 2024) still typically assess abilities in visual perception, recognizing objects, shapes, and colors, and understanding human-centric actions.
 In this paper, we aim to build a benchmark that evaluates AI models’ comprehensive physical reasoning abilities.
@@
 First, this encompasses rigid bodies, ropes, cloths, plastoelastic balls, fluids, and articulated bodies, as well as their couplings, expanding our dataset diversity.
-For instance, Fig 1 (c) portrays the coupling of deformable ropes and articulated pulleys.
-Second, a key feature that distinguishes the proposed ContPhy dataset from existing benchmarks like CLEVRER (Yi et al., 2020) and Physion (Bear et al., ), is the inclusion of varying physical properties and commonsense concepts that are generally acknowledged by common people.
+For instance, Figure 1 (c) portrays the coupling of deformable ropes and articulated pulleys.
+Second, a key feature that distinguishes the proposed ContPhy dataset from existing benchmarks like CLEVRER (Yi et al., 2020) and Physion (Bear et al., 2021), is the inclusion of varying physical properties and commonsense concepts that are generally acknowledged by common people.
 These properties could scarcely be inferred through static images, involving mass, density, tension, friction, stretchiness, bending stiffness, elasticity, and plasticity.
 Such variation can alter the dynamics and generate differing future states, enriching the dataset with a wider range of observable physical behaviors.
-As an example, the liquid’s position in Fig.
-1 (a) is jointly determined by its density and its interaction with the nearby liquids.
+As an example, the liquid’s position in Figure 1 (a) is jointly determined by its density and its interaction with the nearby liquids.
 Scenarios.
-We design four physical dynamic scenarios to comprehensively benchmark models’ cognitive ability on the continuum.
-a) Liquid Dynamics: In Fig.
-2 A, we present a liquid hourglass-like device.
+We design four physical dynamic scenarios to comprehensively benchmark models’ cognitive ability on the continuum, as shown in Figure 2.
+a) Liquid Dynamics.
+In Figure 2 A, we present a liquid hourglass-like device.
 Different liquids, each with unique colors and densities, are released from upper emitters, flowing through fixed sticks, altering directions, and finally reaching containers at the bottom.
@@
 b) Cloths Manipulation.
-As shown in Fig.
-2 B, two cloth pieces with distinct stretching, bending, and frictional characteristics are pulled over objects, inducing potential collision events.
+As shown in Figure 2 B, two cloth pieces with distinct stretching, bending, and frictional characteristics are pulled over objects, inducing potential collision events.
 The released fabric obstructs object views but outlines their shapes through deformations.
@@
 This test assesses models’ ability to discern fabric properties and predict spatial behaviors of concealed objects based on the dynamic 3D surface geometry of the fabric.
-c) Rope Pulley System: In Fig.
-2 C, a wall-mounted arrangement of pulleys, both movable and fixed and anchor points are depicted.
+c) Rope Pulley System.
+In Figure 2 C, a wall-mounted arrangement of pulleys, both movable and fixed and anchor points are depicted.
 Objects with varying masses interact, resulting in diverse motion patterns.
@@
 Additionally, the model is expected to infer numerical relationships between loads’ masses and rope segment tensions.
-d) Soft Ball Dynamics: Fig.
-2 D illustrates a playground with colored obstacles and randomly placed pits.
+d) Soft Ball Dynamics.
+Figure 2 D illustrates a playground with colored obstacles and randomly placed pits.
 Plastoelastic balls with different deformation resistance and yield stress are launched from varying positions, undergoing dynamic movements, including bouncing and permanent deformation.
@@
 We categorize questions into two major groups: Physical Property Questions and Dynamics Questions.
-Fig.
-2 shows the question types of the four scenarios.
-Sample templates are provided in the supplementary material.
+Figure 2 shows the question types of the four scenarios.
+Sample templates are provided in Table 5, 6, 7, and 8 in the Appendix.
 Physical Property Questions.
@@
 We used the Unity engine (Haas, 2014), an efficient platform, to simulate and render videos.
-We follow a bottom-up approach to generate videos and their annotation, involving the following sequential steps: a) Sampling: randomly select scene layouts, camera parameters, and initial conditions to create a diverse set of scenarios.
-b) Initialization: place and configure objects within the scene.
-c) Pre-simulation: conduct a preliminary simulation to evaluate whether the obtained simulation results align with the expected data distribution.
-d) Rendering: generate high-quality videos by configured cameras.
-e) Post-simulation: carry out multiple simulations under varying conditions and record the simulation outputs.
-f) Output: produce rich sensor data and annotation information, encompassing original video, segmentation, bounding boxes, particles, meshes, collision events, configurations, and other simulation raw data required for question generation.
-Supplementary material provides more details.
+We follow a bottom-up approach to generate videos and their annotation, involving the following sequential steps:
+a) Sampling. Randomly select scene layouts, camera parameters, and initial conditions to create a diverse set of scenarios.
+b) Initialization. Place and configure all objects within the scenarios.
+c) Pre-simulation. Conduct a preliminary simulation to evaluate whether the obtained simulation results align with the expected data distribution.
+d) Rendering. Generate high-quality videos with configured cameras.
+e) Post-simulation. Carry out multiple simulations under varying conditions and record the simulation outputs.
+f) Output.
+Produce rich sensor data and annotation information, encompassing original video, segmentation, bounding boxes, particles, meshes, collision events, configurations, and other simulation raw data required for question generation.
+We will provide more details in Section 8.1 in the Appendix.
 Question Generation.
-We develop a question engine to generate question-answering pairs step by step.
-a) Template Design: create massive question and option templates.
-b) Sampling: retrieve the simulation results, combine the properties of possible objects with predefined templates, sample questions, and options accordingly, and determine correct answers.
+We develop a question engine to generate question-answering pairs step by step:
+a) Template Design. Create massive question and option templates.
+b) Sampling.
+Retrieve the simulation results, combine the properties of possible objects with predefined templates, sample questions, and options accordingly, and determine correct answers.
 Target objects possess unique names described by visual attributes like color, shape, orientation, and mobility.
-c) Re-Sampling: ensure a balanced distribution of answers among the options to prevent answer bias.
+c) Re-Sampling. Ensure a balanced distribution of answers among the options to prevent answer bias.
 Question Statistics.
@@
 Across the entire dataset, 20% of questions are counterfactual, 11% are goal-driven, 22% are predictive, and the remaining 46% are related to various physical property questions.
-Further details on the distribution of each question type and templates for each scenario can be found in the supplementary material.
+Further details on the distribution of each question type and templates for each scenario can be found in Section 8.2 in the Appendix.
 4 ContPRO
 Inspired by prior work (Yi et al., 2018, 2020), we propose an oracle neural-symbolic framework named ContPRO for ContPhy.
-As shown in Fig.
-3, we decompose the question-answering task into four main modules, video perception, physical simulation, program parser, and symbolic execution.
+As shown in Figure 3, we decompose the question-answering task into four main modules, video perception, physical simulation, program parser, and symbolic execution.
 Given a raw video, the video perception module detects the objects and their associated static attributes with MASK-RCNN detector (He et al., 2017).
@@
 Physical Simulation.
-We choose DPI-Net (Li et al., 2018) for dynamic prediction for rope and cloth as it has shown reasonable dynamic prediction abilities in different materials (Chen et al., 2022; Bear et al., ).
+We choose DPI-Net (Li et al., 2018) for dynamic prediction for rope and cloth as it has shown reasonable dynamic prediction abilities in different materials (Chen et al., 2022; Bear et al., 2021).
 We also observe that the Material Point Method (MPM) (Sulsky et al., 1995) can estimate physical properties and dynamics better for fluids and objects of varying plasticity with its differentiable mechanism during inference.
@@
 After extracting objects’ static attributes, physical properties, and dynamic trajectories and parsing the natural language query into an executable program, we execute the program with object states as input and output the predicted answer.
-We provide more model details in the supplementary material.
+We provide more model details in Section 10.2 in the Appendix.
 5 Experiments
-Subset | Settings | Blind Models | Visual Models | MLLMs | ContPRO | Human | ||||||
-RND | FRQ | B-LSTM | C-LSTM | MAC | HCRN | ALPRO | Gemini | GPT-4V | ||||
-Rope | Prop. | 30.0 | 53.3 | 54.7 | 52.7 | 53.3 | 51.7 | 60.7 | 35.5 | 48.0 | 71.4 | 84.7 |
-C-Opt. | 51.3 | 51.6 | 74.0 | 74.0 | 74.2 | 74.3 | 76.2 | 48.2 | 42.0 | 75.6 | 90.2 | |
-C-Ques. | 14.7 | 19.0 | 46.0 | 45.0 | 39.8 | 48.1 | 50.7 | 12.0 | 11.3 | 48.8 | 75.0 | |
-G-Opt. | 55.2 | 49.7 | 47.4 | 51.2 | 50.3 | 56.0 | 46.2 | 51.6 | 57.0 | 55.8 | 91.9 | |
-G-Ques. | 4.5 | 11.2 | 7.9 | 6.7 | 6.7 | 2.3 | 1.1 | 10.3 | 12.1 | 2.3 | 84.0 | |
-Fluid | Prop. | 33.3 | 52.7 | 49.3 | 54.0 | 30.0 | 52.7 | 48.0 | 10.0 | 25.0 | 78.0 | 75.8 |
-C-Opt. | 52.9 | 57.9 | 56.1 | 55.0 | 56.5 | 52.6 | 56.8 | 47.3 | 53.3 | 75.7 | 82.5 | |
-C-Ques. | 6.0 | 17.2 | 7.8 | 8.6 | 6.9 | 4.3 | 6.0 | 5.1 | 5.1 | 36.2 | 60.6 | |
-G-Opt. | 59.9 | 63.1 | 57.3 | 57.3 | 51.2 | 67.7 | 62.7 | 44.4 | 53.8 | 77.3 | 75.0 | |
-G-Ques. | 7.5 | 36.3 | 22.5 | 22.5 | 17.5 | 41.3 | 32.5 | 11.3 | 7.5 | 60.0 | 64.3 | |
-P-Opt. | 53.8 | 50.1 | 51.4 | 51.4 | 53.5 | 50.6 | 53.8 | 52.4 | 50.0 | 90.1 | 73.9 | |
-P-Ques. | 4.8 | 12.5 | 12.5 | 12.5 | 12.5 | 1.9 | 12.7 | 5.8 | 13.0 | 68.3 | 42.9 | |
-Cloth | Prop. | 46.7 | 41.3 | 56.7 | 46.7 | 59.3 | 52.0 | 48.0 | 42.0 | 49.0 | 60.0 | 81.4 |
-P-Opt. | 52.2 | 61.7 | 55.2 | 67.5 | 57.9 | 62.0 | 68.8 | 50.1 | 53.0 | 64.7 | 79.6 | |
-P-Ques. | 46.0 | 56.7 | 42.3 | 57.3 | 50.7 | 56.3 | 57.3 | 43.0 | 47.5 | 60.0 | 77.3 | |
-Ball | Prop. | 53.5 | 52.0 | 45.3 | 54.7 | 48.0 | 43.3 | 48.0 | 54.0 | 45.0 | 54.0 | 76.9 |
-C-Opt. | 53.6 | 65.8 | 66.7 | 64.2 | 66.1 | 65.3 | 63.9 | 60.9 | 66.7 | 71.6 | 93.9 | |
-C-Ques. | 30.4 | 48.7 | 43.4 | 41.8 | 3.3 | 28.7 | 40.2 | 29.6 | 46.9 | 57.4 | 90.9 | |
-G-Opt. | 55.9 | 52.1 | 53.3 | 54.1 | 58.1 | 57.0 | 56.3 | 54.1 | 51.4 | 68.1 | 89.7 | |
-G-Ques. | 30.2 | 38.5 | 16.7 | 20.0 | 18.9 | 38.9 | 4.4 | 24.6 | 18.0 | 52.2 | 84.6 | |
-P-Opt. | 50.6 | 67.8 | 68.9 | 67.4 | 64.4 | 61.7 | 65.2 | 51.7 | 45.4 | 92.4 | 72.5 | |
-P-Ques. | 25.9 | 51.7 | 45.5 | 45.5 | 46.6 | 1.1 | 3.4 | 25.9 | 17.2 | 88.6 | 58.8 |
-In this section, we evaluate the ContPhy dataset. We first introduce the experimental setup and then analyze different models’ performance.
+| Subset | Settings | Blind Models | Visual Models | Physical Models | MLLMs | ContPRO | Human | ||||||||
+| RND | FRQ | B-LSTM | C-LSTM | MAC | HCRN | ALPRO | Violet | PhyDNet | PIP | Gemini | GPT-4V | ||||
+| Rope | Prop. | 30.0 | 53.3 | 54.7 | 52.7 | 53.3 | 51.7 | 60.7 | 51.7 | 59.0 | 31.5 | 35.5 | 48.0 | 71.4 | 84.7 |
+| C-Opt. | 51.3 | 51.6 | 74.0 | 74.0 | 74.2 | 74.3 | 76.2 | 76.0 | 77.7 | 75.2 | 48.2 | 42.0 | 75.6 | 90.2 | |
+| C-Ques. | 14.7 | 19.0 | 46.0 | 45.0 | 39.8 | 48.1 | 50.7 | 43.1 | 47.9 | 48.3 | 12.0 | 11.3 | 48.8 | 75.0 | |
+| G-Opt. | 55.2 | 49.7 | 47.4 | 51.2 | 50.3 | 56.0 | 46.2 | 55.2 | 54.4 | 50.6 | 51.6 | 57.0 | 55.8 | 91.9 | |
+| G-Ques. | 4.5 | 11.2 | 7.9 | 6.7 | 6.7 | 2.3 | 1.1 | 1.1 | 5.6 | 2.2 | 10.3 | 12.1 | 2.3 | 84.0 | |
+| Fluid | Prop. | 33.3 | 52.7 | 49.3 | 54.0 | 30.0 | 52.7 | 48.0 | 50.9 | 51.3 | 37.0 | 10.0 | 25.0 | 78.0 | 75.8 |
+| C-Opt. | 52.9 | 57.9 | 56.1 | 55.0 | 56.5 | 52.6 | 56.8 | 60.4 | 59.5 | 49.1 | 47.3 | 53.3 | 75.7 | 82.5 | |
+| C-Ques. | 6.0 | 17.2 | 7.8 | 8.6 | 6.9 | 4.3 | 6.0 | 1.7 | 10.3 | 6.0 | 5.1 | 5.1 | 36.2 | 60.6 | |
+| G-Opt. | 59.9 | 63.1 | 57.3 | 57.3 | 51.2 | 67.7 | 62.7 | 67.3 | 55.9 | 67.7 | 44.4 | 53.8 | 77.3 | 75.0 | |
+| G-Ques. | 7.5 | 36.3 | 22.5 | 22.5 | 17.5 | 41.3 | 32.5 | 41.2 | 40.0 | 41.3 | 11.3 | 7.5 | 60.0 | 64.3 | |
+| P-Opt. | 53.8 | 50.1 | 51.4 | 51.4 | 53.5 | 50.6 | 53.8 | 53.2 | 51.7 | 45.5 | 52.4 | 50.0 | 90.1 | 73.9 | |
+| P-Ques. | 4.8 | 12.5 | 12.5 | 12.5 | 12.5 | 1.9 | 12.7 | 3.8 | 4.8 | 3.8 | 5.8 | 13.0 | 68.3 | 42.9 | |
+| Cloth | Prop. | 46.7 | 41.3 | 56.7 | 46.7 | 59.3 | 52.0 | 48.0 | 55.0 | 58.7 | 54.0 | 42.0 | 49.0 | 60.0 | 81.4 |
+| P-Opt. | 52.2 | 61.7 | 55.2 | 67.5 | 57.9 | 62.0 | 68.8 | 68.2 | 63.5 | 61.6 | 50.1 | 53.0 | 64.7 | 79.6 | |
+| P-Ques. | 46.0 | 56.7 | 42.3 | 57.3 | 50.7 | 56.3 | 57.3 | 55.7 | 47.3 | 46.3 | 43.0 | 47.5 | 60.0 | 77.3 | |
+| Ball | Prop. | 53.5 | 52.0 | 45.3 | 54.7 | 48.0 | 43.3 | 48.0 | 48.0 | 52.7 | 54.0 | 54.0 | 45.0 | 54.0 | 76.9 |
+| C-Opt. | 53.6 | 65.8 | 66.7 | 64.2 | 66.1 | 65.3 | 63.9 | 65.6 | 67.2 | 63.7 | 60.9 | 66.7 | 71.6 | 93.9 | |
+| C-Ques. | 30.4 | 48.7 | 43.4 | 41.8 | 3.3 | 28.7 | 40.2 | 41.8 | 44.3 | 24.6 | 29.6 | 46.9 | 57.4 | 90.9 | |
+| G-Opt. | 55.9 | 52.1 | 53.3 | 54.1 | 58.1 | 57.0 | 56.3 | 57.4 | 57.4 | 54.1 | 54.1 | 51.4 | 68.1 | 89.7 | |
+| G-Ques. | 30.2 | 38.5 | 16.7 | 20.0 | 18.9 | 38.9 | 4.4 | 21.1 | 21.1 | 22.2 | 24.6 | 18.0 | 52.2 | 84.6 | |
+| P-Opt. | 50.6 | 67.8 | 68.9 | 67.4 | 64.4 | 61.7 | 65.2 | 64.4 | 67.4 | 62.9 | 51.7 | 45.4 | 92.4 | 72.5 | |
+| P-Ques. | 25.9 | 51.7 | 45.5 | 45.5 | 46.6 | 1.1 | 3.4 | 2.3 | 17.0 | 6.8 | 25.9 | 17.2 | 88.6 | 58.8 | |
+In this section, we evaluate the ContPhy dataset. We first introduce the experimental setup and then analyze the performance of different models.
 5.1 Experimental Setup
@@
 Each dynamic question is treated as a binary classification task for each question-choice pair.
-For dynamic questions, we report the accuracy for each option and per question.
+For dynamic questions, we report the accuracy of per option per question.
 A question is correct only if all choices in this multiple-choice question are correctly answered.
@@
 HCRN (Le et al., 2020) uses conditional relational networks to learn relations hierarchically in the video, as well as the questions.
-MAC (Hudson & Manning, 2018) has competitive results on previous datasets, which uses co-attention mechanism to model both textual and visual information.
+MAC (Hudson & Manning, 2018) has competitive results on previous datasets, which uses a co-attention mechanism to model both textual and visual information.
 ALPRO (Li et al., 2022a) is a popular model pre-trained on video-text corpus and achieved state-of-the-art results on several video-language datasets.
 We fine-tune ALPRO on our dataset based on the official pre-trained checkpoint.
+Physical Models.
+These specialized models are trained for physical reasoning.
+PhyDNet (Le Guen & Thome, 2020) is a two-branch deep architecture, which explicitly disentangles PDE dynamics from unknown complementary information.
+PIP (Duan et al., 2021) utilizes a deep generative model to model mental simulations, in order to predict future physical interactions.
+To evaluate our benchmark on these physical baselines, we first generated object masks based on each question and fed them into models, together with the video features.
+For the open-ended questions, we added a fully connected layer to predict the answer labels with a cross-entropy loss.
+Multimodal Large Language Models (MLLMs).
 5.2 Evaluation of Physical Reasoning
@@
 Despite these advancements, no visual model has yet achieved top accuracy in all scenarios, underscoring the challenge and significance of our ContPhy.
+Performance of Physical Models.
+Physical models are specialized models designed for special physical tasks.
+These models achieve competitive overall performance and excel in some settings, such as PhyDNet on rope’s and ball’s counterfactual, and PIP on fluid’s goal-driven.
+However, these specialized models also have limitations in some scenarios, such as cloth.
+We hypothesize the reasons are that these models are mainly designed for physical reasoning tasks with simple visual primitives, like sphere collision and movement.
+However, our dataset focuses on continuum objects in diverse environments and different question types, which makes it difficult for these models to grasp the physical rules behind the scenarios.
 Performance of Multimodal Large Language Models.
@@
 We also propose an oracle model to demonstrate the potential to combine recent large language models with traditional particle-based dynamic simulation for effective physical reasoning.
+6 Limitations
+Our proposed benchmark, ContPhy, aims to complement existing physical reasoning benchmarks by encompassing diverse physical property inference across various scenarios and predicting corresponding dynamics.
+However, ContPhy still has limitations.
+Language Diversity.
+While the synthesized questions generated by the question engine can effectively test AI models’ physical reasoning capabilities across diverse scenarios involving different objects, the language diversity remains limited.
+The current set of questions relies on a predefined vocabulary, resulting in a gap compared to natural language.
+Scenario Complexity.
+We have carefully designed four distinct scenarios featuring various objects (e.g., solids, ropes, clothes, and fluids).
+However, real-world physical interactions can be considerably more complex, involving additional objects and physical factors not currently included in the dataset.
 7 Conclusion
@@
 We foresee the ContPhy driving progress in AI perception and reasoning, bridging the gap between human and machine intelligence in the physical world.
+Acknowledgement
+This work was supported by DSO grant DSOCO21072. We would also like to thank the computation support from AiMOS, a server cluster for the IBM Research AI Hardware Center.
+Impact Statement
+This paper presents work whose goal is to advance the field of Physical Commonsense Reasoning.
+We believe our work is useful for 1) evaluating the performance of current existing machine learning models for physical reasoning, and 2) facilitating researchers to develop more powerful AI models with physical commonsense.
+There are no potential negative societal consequences of our work that we feel must be specifically highlighted here.